How do you find the slope? I don’t feel like I did it right

1 x 10-2M CHIM-NSL Frost Looy A Bers Law Exerinen x10 M Repert Sheet: Beers Law Lab Partner Class periodTh p Experimental Data Sheet Week Ose Date 1. Color of stock dye solution selected il p Purple ( Re and Blue te 2) 2. Chemieal mames of dye solutien selected (1 pt): 3. Concentration of stoctdye salation ( .60 x lb 4. Molar Mass of the dyes (1 pt 700 5. Fill in the data table belew Wavelength of ne (This number Abserbaace (2 pt. each) tant) (1 pt.) るE-05 ‘ Concentrahon 13 Slope 2 as 15
Answer #1

In the plot of Absorbance vs concentration, we use the Beer-Lambert's Law, given by:

A=cbc (1)

Its graphic is a line in the plane, where A (absorbance) is the dependent variable, c (concentration) is the independent variable, and the terms \varepsilon b are the slope of the line.

The equation (1) has the general form for a linear function with an intercept of zero, and a slope given by the equation:

\varepsilon b = \frac{A_{2}-A_{1}}{c_2-c_1}

For a pair of points

In this case, we will have the slopes:

\varepsilon b_1 = \frac{A_{2}-A_{1}}{c_2-c_1}=\frac{0.974-0.496}{4.00*10^{-3}-2.00*10^{-3}}=23,900

\varepsilon b_2 = \frac{A_{3}-A_{2}}{c_3-c_2}=\frac{0.496-0.269}{2.00*10^{-5}-1.00*10^{-5}}=22,700

We will have an average of:

\bar{\varepsilon b}=(1/2)(23,900+22,700)= 23,300

According to the data given and the units of the constants.